# latex‘h’ float specifier changed to ‘ht’

### v1

@@ Abstract.
In this paper, we consider the impact of the order flow auction (OFA) in the context of the proposer-builder separation (PBS) mechanism through a game-theoretic perspective. The OFA is designed to improve user welfare by redistributing maximal extractable value (MEV) to the users, in which two auctions take place: the order flow auction and the block-building auction. We formulate the OFA as a multiplayer game, and focus our analyses on the case of two competing players (builders). We prove the existence and uniqueness of a Nash equilibrium for the two-player game, and derive a closed-form solution by solving a quartic equation. Our result shows that the builder with a competitive advantage pays a relatively lower cost, leading to centralization in the builder space. In contrast, the proposer’s shares evolve as a martingale process, which implies decentralization in the proposer (or, validator) space. Our analyses rely on various tools from stochastic processes, convex optimization, and polynomial equations. We also conduct numerical studies to corroborate our findings, and explore other features of the OFA under the PBS mechanism.
@@ 1. Introduction
A blockchain is a decentralized, distributed, and tamper-proof digital ledger that tracks and verifies digital transactions securely without the need for a central authority. Its applications span a wide range of industries, including sustainable energies [27], cryptocurrency [13, 15], healthcare [17, 19, 4], and construction industry [14].
To maintain its decentralized structure, blockchain relies on consensus mechanisms, with the two most widely used being Proof-of-Work (PoW) and Proof-of-Stake (PoS). PoW requires substantial computational power to solve cryptographic puzzles, making it highly energy-intensive. In contrast, PoS is more energy-efficient as it selects validators based on their stake rather than computational effort. Due to its energy efficiency and increasing adoption, we focus on the PoS mechanism in this study.
However, decentralization, the core principle of blockchain technology, faces challenges as large validators gain power in the PoS system. (Validators are often also called proposers; here we use both terms interchangeably.) In PoS, validators with larger holdings have a higher chance of validating new blocks and earning more rewards, increasing wealth concentration and excluding smaller participants [3].
@@
A significant portion of validators’ revenue comes from MEV, which involves strategically reordering, inserting, or censoring transactions within a block to maximize profits. This includes activities like DEX arbitrage, sandwich attacks, and liquidations that exploit inefficiencies in the market. Some of these activities depend solely on the blockchain’s state, using on-chain data to extract value. Others require information from external sources, such as off-chain data, to identify opportunities like CEX-DEX arbitrage [9, 25].
MEV is widely considered as one of the greatest threats to decentralization in blockchain networks, favoring the validators with more resources  [7]. Extracting MEV effectively requires significant capital, advanced strategies, and considerable computational power, which most ordinary validators may not have access to. As a result, well-equipped validators can gain an advantage, further centralizing their control within the network [2, 6, 10, 12, 16, 28].
@@
To distribute MEV fairly among the validators and to prevent centralization, Ethereum introduced the PBS mechanism  [10]. Originally, validators were responsible for both proposing new blocks and constructing their contents. PBS separates these responsibilities into two distinct roles: block builders and block proposers.
Block proposers (or validators) validate and propose blocks; block builders become responsible for assembling blocks. Block builders compete to create the most profitable block, and participate in a block-building auction, offering fees (bids) to the proposer. The proposer then receives the block bundle created by the winning builder and their bid. This allows the proposers to collect significant auction revenue without requiring advanced technical expertise. By fostering competition among the builders, this system diminishes the advantage previously held by sophisticated validators over ordinary ones, facilitating a more balanced distribution of MEV among validators [10, 8, 28, 5].
While PBS can mitigate centralization among the validators to some extent, it tends to create centralization within the builder community. For instance, some builders are exceptionally skilled at exploiting arbitrage opportunities, enabling them to capture significant MEV and consistently win block production opportunities. This concentration of expertise and resources among a few builders can lead to a situation where only a small number of entities control a large portion of the block-building market, thus centralizing power within the builder space [2, 12].
@@
OFA is another mechanism aiming to mitigate centralization and redistribute MEV. While PBS involves the builders and the validators, OFA focuses on the interaction between the users and the builders. It seeks to return a portion of MEV back to the users. In this process, users send their orders to a third-party auction, where block builders or MEV-extracting searchers bid for the exclusive rights to execute strategies on these orders. This approach is conceptually analogous to Payment For Order Flow (PFOF) in traditional finance. Gosselin and Chiplunkar [11], and Ventures [25] outline the structure of OFA as follows:
@@
Figure 1. OFA
Figure  1 presents a flowchart illustrating the OFA process. It is important to note that an order included by a block builder is not necessarily guaranteed on-chain inclusion. The order will only be included if the block builder wins the block-building auction. Therefore, the OFA must provide reliable inclusion guarantees and faces the challenge of the double auction problem, which stems from the interaction between the OFA and the block-building auction [11]. If higher rebates are offered to the users, less MEV can be redistributed to the validators, which may affect the timely inclusion of blocks. Thus, it is crucial to find the optimal balance between user rebates and inclusion fees offered to the validators in the winner selection process [11].
@@
    We formulate the OFA as an $M$-player game, where each builder’s decision variable is the amount of MEV they are willing to pay. Under suitable conditions, we establish that the Nash equilibrium can be characterized by the first-order conditions.
@@
    We study in depth the case when $M=2$, i.e., there are two competing builders. Interestingly, solving for the corresponding Nash equilibrium boils down to solving a univariate quartic equation. We show that there exists a unique Nash equilibrium and derive its closed-form solution. Our analysis reveals that when there are two builders, their equilibrium payments do not scale linearly with the MEV they can extract. Instead, the more capable builder pays relatively less, allowing them to earn higher expected revenue, which drives centralization. Our simulation experiments with three players show a similar pattern, with an interesting variation: the most capable player’s advantage is further amplified, while the gap between the second and the least capable players narrows compared to the two-player case.
@@
    We formulate the evolution of validators’ stake shares as a Pólya urn process with a random replacement matrix. Under this framework, we establish that the stake shares follow a martingale process, aligning with our simulation results, which show that the average stake shares remain nearly constant over time. Furthermore, we analyze the long-term behavior of this process. In the absence of consumption factors (i.e., when no costs are incurred due to staking), we characterize the distribution of the limiting stake shares via functional equations.
Literature Review
Our research contributes to the literature on blockchain centralization. Prior studies have explored various factors driving centralization in blockchain ecosystems, both with and without the PBS mechanism. Reward heterogeneity, a direct cause of centralization, arises from skill disparities among block producers in the absence of PBS [2]. It is further influenced by order flow acquisition, along with the resulting MEV extraction and arbitrage opportunities under PBS [6, 12]. In fact, the role of order flow in centralization has been widely recognized. Empirical studies show that private order flow exacerbates disparities in block-building capacity among builders [28], and those with greater access to private order flow have a higher probability of winning the block-buildling auction [26].
Our research takes into account both the builder heterogeneity in skills and knowledge, and the process of order flow acquisition. Relative to previous studies, we further model order flow acquisition as an auction that explicitly captures its relationship with builders’ block-building capacity, while also incorporating the interaction between the order flow auction and the block-building auction, which remains under-explored in previous studies.
Our research also contributes to the literature on the evolution of validators’ stake shares in the Proof-of-Stake (PoS) system. Rosu and Saleh [18], and Tang [20] show that in a standard PoS system, validators’ stake shares evolve as a martingale process. Tang and Yao [22] analyze stake share evolution under a polynomial voting rule. See also Tang [21] for a review. Building on these works, our research further examines a setting in which the validator’s reward for proposing a block is stochastic and dependent on builders’ bids. In addition, we include consumption factors into the system, recognizing that staking in the pool may incur costs, including opportunity costs, locked funds, and other potential expenses.
Organization of the paper: The remainder of the paper is organized as follows. In Section 2, we present the models of the OFA and PBS. In Section 3, we analyze the Nash equilibrium among the builders, and in Section 4, we study the evolution of stake shares of the validators. We present numerical findings in Section 5. Finally, we conclude with Section 6.
@@ 2. The OFA and PBS Models
  • $a=\mathcal{O}(b)$ means $\frac{a}{b}$ is bounded from above as $b\to\infty$; $a=\Theta(b)$ means $\frac{a}{b}$ is bounded from below and above as $b\to\infty$; and $a=o\left(b\right)$ or $b\gg a$ means $\frac{a}{b}$ decays towards zero as $b\to\infty$.
@@
Builders in a decentralized system are responsible for assembling transaction blocks. They participate in two auctions: the order flow auction and the block-building auction. In the order flow auction, builders compete to acquire users’ order flow, which can provide additional MEV opportunities. By strategically integrating the auctioned order with their existing order flow, they can optimize execution for increased profitability. In the block-building auction, builders bid for the right to propose their assembled block for inclusion on-chain, thereby capturing MEV generated from transaction sequencing and execution.
@@
In our model, we do not distinguish between MEV searchers and builders, but instead treat them as a collective entity. This is motivated by the fact that searchers often direct their order flow to dominant builders or restrict it exclusively to vertically integrated builder-searcher entities [12, 24, 23, 28]. This close coordination, coupled with their shared objective of maximizing MEV, has convinced us to ignore any distinction between their roles, and to focus instead on their combined strategic behavior in the context of MEV extraction.
Let $M\in\mathbb{N}_{+}$ be the total number of builders, and $N\in\mathbb{N}_{+}$ be the total number of validators, which will stay fixed throughout the paper; and let $\left[M\right]$ and $\left[N\right]$ denote the sets of all builders and all validators, respectively. Time is modeled as a discrete sequence indexed by $t=0,1,2,\ldots$. In each period $t$, a single round of the order flow auction and a single round of the block-building auction take place.
@@ 2.1. Builder’s Game
Let $f_{i,t}$ represent the amount of Maximal Extractable Value (MEV) that builder $i$ can capture at round $t$. Assume that $f_{i,t}$ follows a distribution $D_{i,t}$ for each $i\in\left[M\right]$ and for each $t=0,1,2,\ldots$. The expected value of $f_{i,t}$ is given by $\mathbf{E}[f_{i,t}]=\bar{f}_{i,t}$. Suppose builder $i$ has value $v_{i,t}$ for winning the transaction right in the order flow auction at round $t$. Assume that $v_{i,t}$ follows some distribution $F_{i,t}$, for $i\in\left[M\right]$ and $t=0,1,2,\ldots$. The expected value of $v_{i,t}$ is given by $\mathbf{E}[v_{i,t}]=\bar{v}_{i,t}$.
Let $h_{{i,t}}$ denote the total amount that builder $i$ is willing to pay to both users and validators at round $t$. Suppose a fraction $\mu h_{{i,t}}$ is ultimately paid to users, and $(1-\mu)h_{i,t}$ is paid to the selected validator, with $\mu$ determined by the auction design $\left(0\leq\mu\leq 1\right)$. We make the following assumptions.
@@
Builders’ abilities to extract MEV are time-homogeneous.
Assumption 2.3.
The winners of the order flow auction and the block-building auction are determined independently.
@@
We view builders’ capacities to optimize MEV extraction as intrinsic characteristics that remain relatively stable over time. This perspective is in a similar spirit to that of Bahrani et al. [2]. We will omit the time index in the following discussions of builders.
In each auction, the winner is selected randomly, with builder $i$ winning the order flow auction with probability $\frac{h_{i}}{\sum_{j=1}^{M}h_{j}}$ and winning the block-building auction with the same probability. Let $V_{i}$ denote the event that builder $i$ wins the order flow auction, and $Z_{i}$ denote the event that builder $i$ wins the block-building auction:
$\mathbf{I}\left(V_{i}\right)$ =$=\begin{cases}1&\text{if builder $i$ wins the order flow auction,}\\ 0&\text{otherwise.}\end{cases}$
$\mathbf{I}\left(Z_{i}\right)$ =$=\begin{cases}1&\text{if builder $i$ wins the block-building auction,}\\ 0&\text{otherwise.}\end{cases}$
Table 1 presents the four possible outcomes of the two auctions for builder $i$, along with the corresponding revenue (utility) in each case. (Revenue can be regarded as utility, and we use both terms interchangeably in this paper.) Let $h_{-i}$ denote the $M-1$ strategies of all the builders except $i$. The expected utility of builder $i$ is given by:
$\pi_{i}(h_{i}|h_{-i})$ $=\mathbf{E}\left[\mathbf{I}\left(Z_{i}\text{}\right)(f_{i}+v_{i}\mathbf{I}\left(V_{i}\right))-\mu h_{i}\mathbf{I}\left(V_{i}\right)-(1-\mu)h_{i}\mathbf{I}\left(Z_{i}\right)\right],$
@@
Each builder $i$ chooses a bid $h_{i}$ from the strategy space $B_{i}\coloneqq(0,\infty)$ to maximize their expected utility. Let $h_{i}^{*}$ denote the optimal bid, i.e.,
In contrast to  Capponi et al. [6], which introduce an additional player, the order flow provider, and model the order flow acquisition process using a quadratic function, our framework takes a different perspective. Specifically, we model the payment as a proportion of the MEV obtained and introduce randomness in the winner selection process for both the OFA and PBS. As a result, our objective function involves the term $\frac{h_{i}}{\sum_{j=1}^{M}h_{j}}$, reflecting the competitive interaction among the builders.
@@
Win Win $f_{i}+v_{i}-h_{i}$
Win Lose $-\mu h_{i}$
Lose Win $f_{i}-\left(1-\mu\right)h_{i}$
Lose Lose 00
Table 1. Revenue Outcomes for Builder $i$.
@@
Let $M$ denote the total number of builders. Let $B_{i}$ be the set of all possible strategies for builder $i$, where $i\in\left[M\right]$. Let $h=(h_{i},h_{-i})$ be a strategy profile where $h_{-i}$ denotes the $M-1$ strategies of all the builders except $i$. A Nash equilibrium is a strategy profile $h^{*}=(h_{i}^{*},h_{-i}^{*})$ if

### v2

# Analysis of the Order Flow Auction under Proposer-Builder Separation on Blockchain

@@ Abstract.
We study the impact of the order flow auction (OFA) in the context of the proposer-builder separation (PBS) mechanism in blockchains through a game-theoretic perspective. The OFA is designed to improve user welfare by redistributing maximal extractable value (MEV) to the users, in which two sequential auctions take place: the order flow auction and the block-building auction. We formulate the OFA as a multiplayer game, and establish the existence of a Nash equilibrium, and in the two-player case derive a closed-form solution (and prove its uniqueness) via a quartic equation. Our result shows that the builder with a competitive advantage pays a lower cost, leading to a higher revenue, and adding to centralization in the builder space. In contrast, the proposer’s shares evolve as a martingale process, which implies decentralization in the proposer/validator space. Our analyses rely on various tools from stochastic processes, convex optimization, and polynomial equations. We also conduct numerical studies to corroborate our findings, and to bring out other features of the OFA under the PBS mechanism.
@@ 1. Introduction
A blockchain is a decentralized, distributed, and tamper-proof digital ledger that tracks and verifies digital transactions securely without the need for a central authority. Its applications span a wide range of industries, including sustainable energies [wu2018energy], cryptocurrency [hashemi2020cryptocurrency, LAROIYA2020213], healthcare [ratta2021healthcare, cryptography3010003, bell2018applications], and construction industry [KIM20202561].
To maintain its decentralized structure, a blockchain relies on consensus mechanisms, with the two most widely used being Proof-of-Work (PoW) and Proof-of-Stake (PoS). PoW requires substantial computational power to solve cryptographic puzzles, making it highly energy-intensive. In contrast, PoS is more energy-efficient as it selects validators based on their stakes rather than computational power. In view of its energy efficiency and increasing adoption, we focus on the PoS mechanism in this study.
The POS system, however, also faces its own challenges when large validators (or proposers – here we use both terms interchangeably) gain enough power to negatively impact decentralization, the core principle of blockchain technology. This is because validators with larger holdings have a higher chance of validating new blocks and earn more rewards, increasing wealth concentration and excluding smaller participants [bains2022blockchain].
@@
A significant portion of validators’ revenue comes from MEV, which involves strategically reordering, inserting, or censoring transactions within a block to maximize profits. This includes activities like DEX arbitrage, sandwich attacks, and liquidations that exploit inefficiencies in the market. Some of these activities depend solely on the blockchain’s state, using on-chain data to extract value. Others require information from external sources, such as off-chain data, to identify opportunities like CEX-DEX arbitrage [ethereum_mev, monoceros2023].
MEV is widely considered as one of the greatest threats to decentralization in blockchain networks, favoring the validators with more resources  [daian2020]. Extracting MEV effectively requires significant capital, advanced strategies, and considerable computational power, which most ordinary validators may not have access to. As a result, well-equipped validators can gain an advantage, further centralizing their control within the network [bahrani2024, capponi2024, ethereum_pbs, gupta2023, pai2023, yang2024].
@@
To distribute MEV fairly among the validators and to prevent centralization, Ethereum has introduced the PBS mechanism  [ethereum_pbs]. Originally, validators were responsible for both proposing new blocks and constructing their contents. PBS separates these responsibilities into two distinct roles: block builders and block proposers.
Block proposers (or validators) validate/propose blocks; whereas block builders are responsible for assembling blocks. Block builders compete to create the most profitable block, and participate in a block-building auction, offering fees (bids) to the proposer. The proposer then receives the block bundle created by the winning builder and their bid. This allows the proposers to collect significant auction revenue without requiring advanced technical expertise. By fostering competition among the builders, this system diminishes the advantage previously held by sophisticated validators over ordinary ones, facilitating a more balanced distribution of MEV among validators [ethereum_pbs, flashbots2024, yang2024, buterin2021].
While PBS can mitigate centralization among the validators to some extent, it tends to create centralization within the builder community. For instance, some builders are exceptionally skilled at exploiting arbitrage opportunities, enabling them to capture significant MEV and consistently win block-building opportunities. This concentration of expertise and resources among a few builders can lead to a situation where a small number of builders control a large portion of the block-building market, resulting in a concentration of power within the builder space [bahrani2024, gupta2023].
@@
OFA is another mechanism aiming to mitigate centralization and redistribute MEV. While PBS involves the builders and the validators, OFA focuses on the interaction between the users and the builders. It seeks to return a portion of MEV back to the users. In this process, users send their orders to a third-party auction, where block builders or MEV-extracting searchers bid for the exclusive rights to execute strategies on these orders. This approach is conceptually analogous to Payment For Order Flow (PFOF) in traditional finance. gosselin2023, and monoceros2023 outline the structure of OFA as follows:
@@
Figure 1. OFA.
Figure  1 presents a flowchart illustrating the OFA process. It is important to note that an order included by a block builder is not necessarily guaranteed on-chain inclusion. The order will only be included if the block builder wins the block-building auction. Therefore, the OFA must provide reliable inclusion guarantees and faces the challenge of the double auction problem, which stems from the interaction between the OFA and the block-building auction [gosselin2023]. If higher rebates are offered to users, less MEV can be redistributed to validators, which may affect the timely inclusion of blocks. Thus, it is crucial to find the optimal balance between user rebates and inclusion fees offered to validators in the winner selection process [gosselin2023].
Interplay Between PBS and OFA
While PBS and OFA target different parts of the blockchain system, with PBS redistributing MEV between builders and validators and OFA seeking to return MEV to users, they are interconnected. In practice, these two mechanisms operate sequentially: builders first acquire users’ order flow through OFA, then aggregate this order flow with other available transactions to construct a block, which they submit for inclusion via PBS. This introduces a dependency: the value that a builder can extract from the order flow may influence their bid in PBS. Conversely, the likelihood of winning the block-building auction in PBS constrains the amount a builder is willing to pay for order flow in OFA. Hence, the two auctions form a coupled system in which builders optimize their total bid across both mechanisms.
@@
    We formulate the OFA as an $M$-player game, where each builder’s decision variable is the amount of MEV they are willing to pay. Under suitable conditions, we establish the existence of a Nash equilibrium and provide its characterization.
@@
    We study in depth the case when $M=2$, i.e., there are two competing builders. Interestingly, solving for the corresponding Nash equilibrium boils down to solving a univariate quartic equation. We show that there exists a unique Nash equilibrium and derive its closed-form solution. Our analysis reveals that when there are two builders, their equilibrium payments do not scale linearly with the MEV they can extract. Instead, the more capable builder pays less, and hence earns a higher expected revenue, which drives centralization. Our simulation experiments with three players show a similar pattern, with an interesting variation: the most capable player’s advantage is further amplified, while the gap between the second and the least capable players narrows compared to the two-player case.
@@
    We formulate the evolution of validators’ stake shares as a Pólya urn process with a random replacement matrix. With this model, we show the stake shares follow a martingale process. This aligns with our simulation results, which demonstrate that the average stake shares remain nearly constant over time. Furthermore, we analyze the long-term behavior of this process. In the absence of consumption factors (i.e., when no costs are incurred due to staking), we characterize the distribution of the limiting stake shares via functional equations.
Related Literature
Our paper relates closely to the literature on centralization in blockchains. Prior studies have explored various factors driving centralization, both with and without the PBS mechanism. Reward heterogeneity, a direct cause of centralization, arises from skill disparities among block producers in the absence of PBS [bahrani2024]. It is further influenced by order flow acquisition, along with the resulting MEV extraction and arbitrage opportunities under PBS [capponi2024, gupta2023]. In fact, the role of order flow in centralization has been widely recognized. Empirical studies show that private order flow exacerbates disparities in block-building capacity among builders [yang2024], and those with greater access to private order flow have a higher probability of winning the block-buildling auction [wang2024privateorderflowsbuilder].
Our study takes into account both the builder heterogeneity in skills and knowledge, and the process of order flow acquisition. Relative to previous studies, we further model order flow acquisition as an auction that explicitly captures its relationship with builders’ block-building capacity, while also incorporating the interaction between the order flow auction and the block-building auction, which appears under-explored in prior works.
Our research also contributes to the literature on the evolution of validators’ stake shares in the Proof-of-Stake (PoS) system. rosu2021market, and Tang22 show that in a standard PoS system, validators’ stake shares evolve as a martingale process. tang2024polynomialvotingrules analyze stake share evolution under a polynomial voting rule. See also Tang24 for a review. Building on these works, our research further examines a setting in which the validator’s reward for proposing a block is stochastic and dependent on builders’ bids. In addition, we include consumption factors into the system, recognizing that staking in the pool may incur costs, including opportunity costs, locked funds, and other potential expenses.
Organization of the paper: The remainder of the paper is organized as follows. In Section 2 we introduce the two models for OFA and for PBS. We analyze the Nash equilibrium among the builders in Section 3, and then study the evolution of stake shares of the validators in Section 4. Numerical examples and findings are presented in Section 5, and concluding remarks in Section 6.
@@ 2. The OFA and PBS Models
  • $a=\mathcal{O}(b)$ means $\frac{a}{b}$ is bounded from above as $b\to\infty$, and $a\sim b$ means $\frac{a}{b}$ converges to $1$ as $b\to\infty$.
@@
Builders in a decentralized system are responsible for assembling transaction blocks. They participate in two sequential auctions: the order flow auction and the block-building auction. In the order flow auction, builders compete to acquire users’ order flow, which can provide additional MEV opportunities. By strategically integrating the auctioned order with their existing order flow, they can optimize execution for increased profitability. In the subsequent block-building auction, builders bid for the right to propose their assembled block for inclusion on-chain, thereby capturing MEV generated from transaction sequencing and execution.
@@
In our model, we do not distinguish between MEV searchers and builders, but instead treat them as a collective entity. This is motivated by the fact that searchers often direct their order flow to dominant builders or restrict it exclusively to vertically integrated builder-searcher entities [gupta2023, titan2023builder, ethresear2024buildersbehavioralprofiles, yang2024]. This close coordination, coupled with their shared objective of maximizing MEV, has convinced us to ignore any distinction between their roles, and to focus instead on their combined strategic behavior in the context of MEV extraction.
Let $M\in\mathbb{N}_{+}$ be the total number of builders, and $N\in\mathbb{N}_{+}$ be the total number of validators, both of which remain fixed throughout the paper; and let $\left[M\right]$ and $\left[N\right]$ denote the sets of all builders and all validators, respectively. Builders first participate in a single round of the order flow auction, followed by a single round of the block-building auction.
@@ 2.1. Builder’s Game
Let $f_{i}$ represent the amount of Maximal Extractable Value (MEV) that builder $i$ can capture. Assume that $f_{i}$ follows a distribution $D_{i}$ for each $i\in\left[M\right]$. The expected value of $f_{i}$ is given by $\mathbf{E}[f_{i}]=\bar{f}_{i}$. Suppose builder $i$ has value $v_{i}$ for winning the transaction right in the order flow auction. Assume that $v_{i}$ follows some distribution $F_{i}$, for $i\in\left[M\right]$. The expected value of $v_{i}$ is given by $\mathbf{E}[v_{i}]=\bar{v}_{i}$.
Let $r_{{i}}$ denote the total amount that builder $i$ bids to both users and the selected validator. $r_{i}$ is modeled as a random variable parametrized by $h_{i}$, such that $\mathbb{E}(r_{i})=h_{i}$, with support $\left[r_{i}^{\min},r_{i}^{\max}\right]$, where $r_{i}^{\min}>0$. By definition, it follows that $h_{i}\in\left[r_{i}^{\min},r_{i}^{\max}\right]$. Suppose a fraction $\mu r_{{i}}$ is ultimately bid to users, and $(1-\mu)r_{i}$ is bid to the selected validator, with $\mu$ determined by the design of the order flow auction $\left(0\leq\mu\leq 1\right)$. Note that the two auctions are sequential, and that builders must report, in the first auction—the order flow auction—the total amount they intend to bid across both the OFA and the PBS auctions. A fraction of this reported amount is bid to users through the OFA, while the remainder is bid to the selected validator through PBS [gosselin2023]. We make the following assumption.
@@
The winners of the order flow auction and the block-building auction are determined independently, and both of them are first-price auctions. The probability that $r_{i}$ is the highest among $r_{1},\dots,r_{M}$ is given by
$\mathbf{P}\left(r_{\left(M\right)}=r_{i}\right)=\frac{h_{i}}{\sum_{j=1}^{M}h_{j}},$ (2.1)
where $r_{\left(M\right)}=\max\left\{r_{1},r_{2},\dots,r_{M}\right\}$.
@@
Equation 2.1 is based on the Plackett-Luce model from the theory of social choice. In this framework, each player $i$ is assigned a positive parameter $h_{i}$ representing their latent “ability” or “competitiveness,” which may reflect heterogeneity among builders, such as differences in resources or technical capabilities. The model implies that the probability of player $i$ submitting the highest bid $r_{i}$ among all players is proportional to $h_{i}$, normalized by the sum $\sum_{j=1}^{M}h_{j}$.
Let $V_{i}$ denote the event that builder $i$ wins the order flow auction, and $Z_{i}$ denote the event that builder $i$ wins the block-building auction:
$\mathbf{I}\left(V_{i}\right)$ $=\begin{cases}1&\text{if builder $i$ wins the order flow auction,}\\ 0&\text{otherwise.}\end{cases}$
$\mathbf{I}\left(Z_{i}\right)$ $=\begin{cases}1&\text{if builder $i$ wins the block-building auction,}\\ 0&\text{otherwise.}\end{cases}$
Table 1 presents the four possible outcomes of the two auctions for builder $i$, along with the corresponding revenue (utility) in each case. (Revenue can be regarded as utility, and we use both terms interchangeably in this paper.) Let $h_{-i}$ denote the M1M-1 strategies of all the builders except $i$. The expected utility of builder $i$ is given by:
$\pi_{i}(h_{i}|h_{-i})$ $=\mathbf{E}\left[\mathbf{I}\left(Z_{i}\text{}\right)(f_{i}+v_{i}\mathbf{I}\left(V_{i}\right))-\mu r_{i}\mathbf{I}\left(V_{i}\right)-(1-\mu)r_{i}\mathbf{I}\left(Z_{i}\right)\right],$
@@
Each builder $i$ chooses a bid with expectation $h_{i}$ from the strategy space $B_{i}\coloneqq\left[r_{i}^{\min},r_{i}^{\max}\right]$, where $r_{i}^{\min}>0$, to maximize their expected utility. Let $h_{i}^{*}$ denote the optimal bid, i.e.,
$h_{i}^{*}=\arg\max_{h_{i}\in B_{i}}\bar{f}_{i}\frac{h_{i}}{\sum_{j=1}^{M}h_{j}}+\bar{v}_{i}\left(\frac{h_{i}}{\sum_{j=1}^{M}h_{j}}\right)^{2}-\frac{h_{i}^{2}}{\sum_{j=1}^{M}h_{j}}.$
In contrast to  capponi2024, which introduce an additional player, the order flow provider, and model the order flow acquisition process using a quadratic function, our framework takes a different perspective. Specifically, we model the payment as a proportion of the MEV obtained and introduce randomness in the winner selection process for both the OFA and PBS. As a result, our objective function involves the term $\frac{h_{i}}{\sum_{j=1}^{M}h_{j}}$, reflecting the competitive interaction among the builders.
@@
Win Win $f_{i}+v_{i}-r_{i}$
Win Lose $\mu r_{i}$
Lose Win $f_{i}-\left(1-\mu\right)r_{i}$
Lose Lose 0
Table 1. Revenue Outcomes for Builder ii.
@@
Let $M$ denote the total number of builders. Let $B_{i}$ be the set of all possible strategies for builder $i$, where $i\in\left[M\right]$. Let $h=(h_{i},h_{-i})$ be a strategy profile where $h_{-i}$ denotes the M1M-1 strategies of all the builders except $i$. A Nash equilibrium is a strategy profile $h^{*}=(h_{i}^{*},h_{-i}^{*})$ if